# Meta-Property Graphs: Extending Property Graphs with Metadata Awareness and Reification

### v1

@@ 1 Introduction
Modern data engineering applications demand flexibility and agility more than ever before. For example, during exploratory analytics, one typically starts in a schema-less situation, where structure is discovered over time. Heterogeneity in what is a data value versus what is an attribute name is inherent during discovery, profiling, and exploration. As another example, during integration of data sources, such as when building a knowledge graph, data and schema heterogeneity is the default. One source’s data values correspond to another source’s node labels, and in yet another data source, these correspond to node property names; further, what is a subgraph in one source corresponds to a node in a different data source. Modern data management solutions must fully embrace and support heterogeneity and diversity in the modeling distinctions between data and metadata.
@@
In data modelling, metadata is commonly understood in two distinct ways. The first and most common form is attribute data, which refers to the characteristics or details associated with a piece of information, such as a relational database table’s attributes (i.e., column names) or the table’s name itself. For instance, when storing a person’s contact information, attributes like name, email address, phone number, and address are essential metadata. On the other hand, another common form of metadata is reification. This concept involves aggregating complex relationships or sets into a new entity or object, making it easier to manage and analyze. In the context of entity-relationship modeling, reification is often used to transform relationship sets into entity sets, allowing for more effective data modeling and manipulation. For example, in an online shopping platform, reifying the relationship between customers and orders could result in a new entity representing the order history or customer preferences. [^1] Other more complex types of metadata, such as reflective or active data [7, 9, 18], are beyond the scope of our current discussion.
@@
The ISO standard Property Graph (PGs) model has gained popularity in graph data management and is widely adopted, e.g., in graph DB systems such as Neo4j, Tigergraph, and Amazon Neptune, as well as relational DB systems such as DuckDB which implement the ISO extensions to SQL for PG querying. In a property graph nodes and edges are labeled and also have associated sets of property name/value pairs (e.g., a node labeled Person with property Birthdate having data value 11-11-2001; here Person and Birthdate are attribute metadata). While PGs offer a model closely aligned with conceptual domain representations, the model makes (1) a strict distinction between metadata and data, and (2) has no support for reification. In contrast, the W3C’s Resource Description Framework (RDF) for modeling graph data treats everything as data, including attributes of nodes and edges, and natively supports reification. This capability makes RDF particularly useful for applications requiring sophisticated data validation, verification, querying, and governance of metadata. The lack of metadata awareness in PGs, such as the ability to query and audit metadata, represents a significant limitation compared to RDF’s versatility.
The vision.
@@
In this paper we present a vision for overcoming barriers to flexible management of data-metadata heterogeneity in PG data management applications. Towards this, we introduce Meta-Property Graphs (MPG), a fully backwards compatible extension of the PG model that addresses limitations in representing and querying metadata. Our approach enables first-class treatment of labels and properties as queryable objects, as well as reification of subgraphs. On this foundation, we further propose MetaGPML, an fully backwards compatible extension of the Graph Pattern Matching Language (GPML), the core language at the heart of the ISO standard GQL for PG querying. We give a complete formal specification of MPG and MetaGPML, providing the foundations for our vision. Furthermore, we demonstrate how these contributions facilitate advanced data analytics, integration, and governance tasks that are challenging or impossible with current PG systems.
@@ 2 Related Work
The PG model represents a design point on the continuum between the Relational (fixed metadata) and RDF (no fixed metadata) data models. While by default schema-less, supporting agile modeling and development, PGs nonetheless make a strict distinction between data values and metadata values. Of course, all points along this design continuum are equally important, finding their applications and use-cases. And indeed, data cleaning, wrangling, integration, and exchange are very often about moving (meta)data along the continuum. What has been missing is an appropriate design for PGs to fully meet modeling and querying needs of data management scenarios such as these (see Chapter 2 of [5] for a survey of design approaches for PG modeling).
In the research literature, dealing with the challenges of data-metadata heterogeneity was studied in the context of relational data integration and data integration on the web, leading to solutions for relational and XML data-metadata mapping and exchange (e.g., [4, 10, 15]) and seminal work such as the SchemaSQL and FISQL data-metadata query languages for relational databases [16, 21]. Our work approaches these challenges in the new context of the ISO standards for graph data management.
In another direction, there is a growing body of work on mappings between graph data models [1, 6, 14, 20, 19] and generalizing both PG and RDF [13, 17, 3]. Our vision is orthogonal and complementary to these investigations, aiming at extending the PG model and the standardized GQL and SQL/PGQ query languages with seamless data-metadata functionalities.
@@ 3 A guided tour of working with MPGs
Figure 1 illustrates a simple MPG, demonstrating how the MPG model enables direct querying of metadata as data objects. The sample database represents publications, indexing databases, and persons with their relationships, modeled using four types of data objects: edges ($E_{G}$), nodes ($N_{G}$), properties ($P_{G}$), and labels ($L_{G}$). Note that properties and labels are now first class citizens, unlike in the current PG standard.
@@ 3.1 Working with Data, Labels, and Properties
[…]
@@ 3.2 Working with Reification
MPG and MetaGPML enable reification as the second type of metadata on property graphs. A sub-structure of an MPG can be reified as a node, thereby making it a first class citizen. This flexibility allows for partial relationship reification, which can be useful for privacy preservation. The model also permits a node to reify its own properties alongside other data objects. In Figure 1, we represent the statement “Rose assigned Lee as a reviewer on 5th November 2024” using a node that induces a sub-structure containing node $n_{1}$, relevant properties, and outgoing review edges. Query $Q_{5}$ retrieves the name of the person who assigns a reviewer:
@@ 4 Meta-Property Graph Model
where $E=E_{d}\cup E_{u}$ the set of compatible property sets =$\mathbb{C}=\{C\subseteq P\mid\forall p_{1},p_{2}\in C,p_{1}\neq p_{2}\Rightarrow Key(p_{1})\neq Key(p_{2})\}$, $Key(p)=\pi_{1}(\upsilon(p))$, $Val(p)=\pi_{2}(\upsilon(p))$, and $\forall o_{1},o_{2}\in N\cup E,o_{1}\neq o_{2}\Rightarrow\sigma(o_{1})\cap\sigma(o_{2})=\varnothing$ and $\bigcup_{o\in N\cup E}\sigma(o)=P$.
@@
The constraints we have imposed in our design of MPGs are very minimalistic, ensuring only that all properties are uniquely assigned to edges and nodes, and that reification is well-founded. We have deliberately avoided imposing unwarranted constraints. For instance, there is no constraint preventing a node from having an outgoing edge to a subgraph containing one of the node’s own properties. This flexibility allows for representing scenarios such as “Mary legally changed her first name” or “Mary audits all names in the database (including her own)”. Similarly, there is no constraint requiring that for each edge in $\rho(n)$, the source and target of the edge must also be in $\rho(n)$. This permits representations like “Mary is confident that Bob edited a book, but not confident about which book”, where only the node representing Bob and Bob’s outgoing edge labeled ’edited’ are included, without the target of this edge. Of course, such additional constraints can be added (or existing constraints can be removed) as appropriate to the specific application domain.
@@ 5 Meta-Property Graph Pattern Matching Language
We next formalize our Graph Pattern Matching Language for Metaproperty Graphs (MetaGPML) as an extension to the Graph Pattern Matching Language (GPML) [8, 12, 11]. GPML is the core element of both SQL/PGQ and GQL languages standardized by ISO. By extending GPML we ensure backward compatibility in the sense that every GPML query is a MetaGPML query.
@@ 5.1 Syntax
Let $\mathcal{X}$ be a countable infinite set of variables. We introduce descriptors $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ in MetaGPML, consistent with GPML[12]. For $x,y\in\mathcal{X},\ \ell\in\mathcal{L},\ \mathcal{k}\in\mathcal{K},$ and $\mathcal{v}\in\mathcal{V}$, we have
Descriptors: $\delta_{1}$ $:=x\mid:\ell\mid::\pi\mid x:\ell\mid:\ell::\pi\mid x::\pi\mid x:\ell::\pi$ (1) :?$y\mid x:?y\mid:?y::\pi\mid x:?y::\pi$ $\delta_{2}$ $:=x\mid:\ell\mid:?y\mid x:\ell\mid x:?y$ $\delta_{3}$ $:=x$
Here, $\pi$ is a pattern that describes different object patterns and some operations over them in a meta-property graph or sub-structure.
Patterns: $\begin{array}[]{llr}\pi:=&(\delta_{1})\ \mid\ ()\ \mid\ (\delta_{1}).\delta_{3}\ \mid\ ().\delta_{3}&\text{(node pattern)}\\ &\mid\ -[\delta_{2}]->\ \mid\ <-[\delta_{2}]-\ \mid-[\delta_{2}]-&\text{(edge pattern)}\\ &\mid\ -[]->\ \mid\ <-[]-\ \mid-[]-&\\ &\mid\ -[\delta_{2}].\delta_{3}->\ \mid\ <-[\delta_{2}].\delta_{3}-&\\ &\mid\ -[\delta_{2}].\delta_{3}-&\\ &\mid\ \{\delta_{3}\}\ \mid\ \{\}&\text{(property pattern)}\\ &\mid\ \ \mid\delta_{3}\mid\ \mid\ \mid\mid&\text{(label pattern)}\\ &\mid\ \pi\pi&\text{(concatenation)}\\ &\mid\ \pi+\pi&\text{(union)}\\ &\mid\ \pi\ \text{{WHERE}}\ \Phi&\text{(conditioning)}\\ \Pi:=&\pi\mid\Pi,\Pi&\text{(graph pattern)}\\ \end{array}$ (13)
We define $\Phi$ as the expression for conditions that can be used in a pattern after the WHERE keyword as follows, considering $p\in P$ and $c\in\mathcal{K}\cup\mathcal{V}\cup L$.
Expressions and $\begin{array}[]{llr}\mathcal{E}:=&\ x\mid\ x.\mathcal{k}\ \mid\ \mathcal{v}\ \mid\ \mathcal{k}\ \mid\ \ell&\hbox{\multirowsetup\text{(expressions)}}\\ &\ \mid\ \text{{KEY}}(p)\ \mid\ \text{{VAL}}(p)&\\ \Phi:=&\mathcal{E}=\mathcal{E}\ \mid\ \mathcal{E}<\mathcal{E}\ \mid\ x:\ell&\hbox{\multirowsetup\text{(conditions)}}\\ &\mid\text{{SUBSETEQ}}(x,\ y)\ \mid\ c\ \text{{ELEMENTOF}}\ y&\\ &\mid\ \Phi\ \text{{AND}}\ \Phi\ \mid\ \Phi\ \text{{OR}}\ \Phi\ \mid\ \text{{NOT}}\ \Phi&\end{array}$ (19)
Finally, based on (13), (19), we formulate the definition of clauses and queries.
Clauses and queries: $\begin{array}[]{ll}\mathcal{C}:=&\text{{MATCH}}\ \Pi\\ &\ \mid\ \text{{FILTER}}\ \Phi\\ \mathcal{Q}:=&\mathcal{C}\mathcal{Q}\\ &\mid\ \text{{RETURN}}\ \mathcal{E}_{1}\ \text{{AS}}\ x_{1},\ ...\ ,\mathcal{E}_{n}\ \text{{AS}}\ x_{n}\\ \end{array}$ (24)
@@ 5.2 Semantics
We next present the complete semantics of our language, as a clear unambiguous foundation for further study and implementation.
@@
(Bindings) A binding $\beta:\mathcal{X}\rightarrow\mathbb{V}$ assigns variables $x\in\mathcal{X}$ to values $v\in\mathbb{V}$, where $\mathbb{V}=\mathcal{I}\cup\mathcal{L}\cup\mathcal{K}\cup\mathcal{V}$. We denote this as $(x_{1}\mapsto v_{1},...,x_{n}\mapsto v_{n})$, where $x_{1},...,x_{n}$ are the variables in the domain of $\beta$ $(\text{Dom}(\beta))$ and $v_{1},...,v_{n}$ are their corresponding values. We denote the empty binding as ()()\,( ).
@@
Node pattern matching: $\llbracket()\rrbracket_{G}=\{()\mid n\in N\}$ (25) $\llbracket(x)\rrbracket_{G}=\{(x\mapsto n)\mid n\in N\}$ (26) $\llbracket(x:\ell)\rrbracket_{G}=\{(x\mapsto n)\mid n\in N,\ell\in\mu(\lambda(n))\}$ (27) $\llbracket(x:?y)\rrbracket_{G}=\{(x\mapsto n,y\mapsto l)\mid n\in N,l=\lambda(n)\}$ (28) $\llbracket(x).z\rrbracket_{G}=\{(x\mapsto n,z\mapsto p)\mid n\in N,p\in\sigma(n)\}$ (29)
@@
Meta-nodes are a key concept in MetaGPML, defined as nodes to which diverse elements are assigned based on the concept of sub-structure. This feature enables reification within the Metaproperty Graph, allowing for a more nuanced representation of complex relationships and structures.
@@
Meta-node pattern matching: $\llbracket(x:\ell::\pi)\rrbracket_{G}=\llbracket(x:\ell)\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (30)
The meta-node semantic definition involves a full join operation between two patterns divided by :::absent:::: : in the node pattern descriptor. Here, $\llbracket(x:\ell)\rrbracket_{G}$ represents a simple node pattern matching, while $\llbracket\pi\rrbracket_{G_{n}}$ is a pattern matching within the $G_{n}$ sub-structure. We define the semantics of edges in our model, supporting both directed and undirected edges.
Edge pattern matching: $\llbracket-[]->\rrbracket_{G}$ =$=\{()\mid e\in E_{d}\}$ (31) $\llbracket-[x]->\rrbracket_{G}$ =$=\{(x\mapsto e)\mid e\in E_{d}\}$ (32) $\llbracket-[:\ell]->\rrbracket_{G}$ =$=\{()\mid e\in E_{d},\ell\in\mu(\lambda(e))\}$ (33) $\llbracket-[x:\ell]-\rrbracket_{G}$ =$=\{(x\mapsto e)\mid e\in E_{u},\ell\in\mu(\lambda(e))\}$ (34) $\llbracket-[x:?y]->\rrbracket_{G}$ =$=\{(x\mapsto e,y\mapsto l)\mid e\in E_{d},l=\lambda(e)\}$ (35) $\llbracket-[x].z->\rrbracket_{G}$ =$=\{(x\mapsto e,z\mapsto p)\mid e\in E_{d},p\in\sigma(e)\}$ (36)
Equations 31-34 match various edge patterns, including directed and undirected edges with or without labels. Equation 35 allows assigning the edge’s label set to variable $y$, while 36 enables assigning the edge’s properties to variable $z$.
@@
Property and label objects: $\llbracket\{x\}\rrbracket_{G}$ =$=\{(x\mapsto p)\mid p\in P\}$ (37) $\llbracket|x|\rrbracket_{G}$ =$=\{(x\mapsto l)\mid l\in L\}$ (38)
Pattern concatenation, union, and conditioning: $\llbracket\pi_{1}\pi_{2}\rrbracket_{G}$ =$=\{\beta_{1}\Join\beta_{2}\mid\beta_{i}\in\llbracket\pi_{i}\rrbracket_{G},i=1,2\text{ and }\beta_{1}\sim\beta_{2}\}$ (39) $\llbracket\pi_{1}+\pi_{2}\rrbracket_{G}$ =$=\{\beta\cup\beta^{\prime}\mid\beta\in\llbracket\pi_{1}\rrbracket_{G}\cup\llbracket\pi_{2}\rrbracket_{G}\}$ (40) $\llbracket\pi\ \text{{WHERE}}\ \Phi\rrbracket_{G}$ =$=\{\beta\in\llbracket\pi\rrbracket_{G}\mid\llbracket\Phi\rrbracket^{\beta}_{G}=True\}$ (41)
In (40), $\beta^{\prime}$ maps any variable in ${\llbracket\pi_{1}+\pi_{2}\rrbracket}_{G}$ not in $\beta$’s domain to $Null$.
@@
The semantics $\llbracket\mathcal{E}\rrbracket^{\beta}_{G}$ of an expression $\mathcal{E}$ is computed with respect to binding $\beta$ over graph $G$. For a variable $x$ in $\beta$’s domain, $\llbracket x\rrbracket^{\beta}_{G}=\beta(x)$.
Node and property values: $\llbracket x.\mathcal{k}\rrbracket^{\beta}_{G}=\begin{cases}Val(p)&\text{if }\beta(x)\in N\cup E,p\in\sigma(\beta(x)),Key(p)=\mathcal{k}\\ Null&\text{otherwise}\end{cases}$ (43)
Property key and value operations: $\llbracket\text{{KEY}}(p)\rrbracket^{\beta}_{G}$ $=Key(p)\text{ for }p\in P$ (44) $\llbracket\text{{VAL}}(p)\rrbracket^{\beta}_{G}$ $=Val(p)\text{ for }p\in P$ (45)

### v2

@@ 1 Introduction
Modern data engineering applications increasingly demands flexibility and agility in handling data and metadata more than ever before. During exploratory analytics, structure emerges gradually without a predefined schema. Heterogeneity in what is a data value versus what is an attribute name is inherent during discovery, profiling, and exploration. In data sources integration scenarios, such as building knowledge graphs, data and schema heterogeneity is inevitable: what appears as a data value in one source might be a node label or property name in another, while a subgraph in one source might correspond to a single node elsewhere. Modern data management solutions must therefore fully support heterogeneity and fluid boundaries between data and metadata.
@@
Metadata is commonly understood in two distinct ways. The first is attribute metadata- characteristics associated with data objects, such as column names or table names in relational databases or node labels in graphs. For instance, when storing contact information, attributes like name and email are metadata. The second form is reification- representing an aggregation of complex relationships or structures into a new entity or data object, making it easier to manage and analyze. In the context of entity-relationship modeling, reification is often used to transform relationship sets into entity sets, allowing for more effective data modeling and manipulation. For example, in an e-commerce system, reifying the relationship between customers and orders creates an order history entity that can be analyzed and queried. [^1] Other more complex types of metadata, such as reflective or active data [stijn, Fletcher14, divesh], are beyond the scope of our current discussion.
@@
The ISO standard Property Graph (PGs) model has gained popularity in graph data management and is widely adopted, e.g., in graph DB systems such as Neo4j, Tigergraph, and Amazon Neptune, as well as relational DB systems such as DuckDB which implement the ISO extensions to SQL for PG querying. In a property graph nodes and edges are labeled and also have associated sets of property name/value pairs (e.g., a node labeled Person with property Birthdate having data value 11-11-2001; here Person and Birthdate are attribute metadata). While PGs offer a model closely aligned with conceptual domain representations, the model makes (1) a strict distinction between metadata and data, and (2) has no support for reification. In contrast, the W3C’s Resource Description Framework (RDF) for modeling graph data treats everything as data, including attributes of nodes and edges, and natively supports reification. This capability makes RDF particularly useful for applications requiring sophisticated data validation, verification, querying, and governance of metadata. There is an opportunity to bring these powerful features to PGs while preserving their core design principles and inherent strengths in handling data heterogeneity.
@@
In this paper we present a vision for overcoming barriers to flexible management of data-metadata heterogeneity in PG data management applications. Towards this, we introduce Meta-Property Graphs (MPG), a fully backwards compatible extension of the PG model that addresses limitations in representing and querying metadata. Our approach enables first-class treatment of labels and properties as queryable objects, as well as reification of subgraphs. On this foundation, we further propose MetaGPML, a fully backwards compatible extension of the Graph Pattern Matching Language (GPML), the core language at the heart of the ISO standard GQL for PG querying. We give a complete formal specification of MPG and MetaGPML, providing the foundations for our vision. Furthermore, we demonstrate how these contributions facilitate advanced data analytics, integration, and governance tasks that are challenging or impossible with current PG systems.
@@ 2 Related Work
The PG model represents a design point on the continuum between the Relational (fixed metadata) and RDF (no fixed metadata) data models. While by default schema-less, supporting agile modeling and development, PGs nonetheless make a strict distinction between data values and metadata values. Of course, all points along this design continuum are equally important, finding their applications and use-cases. And indeed, data cleaning, wrangling, integration, and exchange are very often about moving (meta)data along the continuum. What has been missing is an appropriate design for PGs to fully meet modeling and querying needs of data management scenarios such as these (see Chapter 2 of [bonifati] for a survey of design approaches for PG modeling).
In the research literature, dealing with the challenges of data-metadata heterogeneity was studied in the context of relational data integration and data integration on the web, leading to solutions for relational and XML data-metadata mapping and exchange (e.g., [BonifatiCHLPC10, fletcher_towards_2009, HernandezPT08]) and seminal work such as the SchemaSQL and FISQL data-metadata query languages for relational databases [lakshmanan_schemasql_2001, wyss_relational_2005]. Our work approaches these challenges in the new context of the ISO standards for graph data management.
In another direction, there is a growing body of work on mappings between graph data models [mappings_4, mappings_1, mappings_2, mappings_3, StoicaFS19] and generalizing both PG and RDF [gelling_bridging_2023, onegraph, multilayer_graphs]. Our vision is orthogonal and complementary to these investigations, aiming at extending the PG model and the standardized GQL and SQL/PGQ query languages with seamless data-metadata functionalities.
@@ 3 A guided tour of working with MPGs
Figure 1 illustrates a simple MPG, demonstrating how the MPG model enables direct querying of metadata as data objects. The sample database represents publications, indexing databases, and persons with their relationships, modeled using four types of data objects: edges ($E_{G}$), nodes ($N_{G}$), properties ($P_{G}$), and labels ($L_{G}$). Note that properties and labels are now first class citizens, unlike in the current PG standard. Table 1 introduces the pattern notation used in our examples, which will be formally defined in Section 5.
Table 1: Data objects pattern notation Nodes
(x:l) node variable (nv.) x with label l
(x:l).z nv. x with label l and property variable z
[…]
Edges
-[x:l]-> edge variable x with label l
-[x:?y]-> edge variable x with label set variable y
-[x].z-> edge variable x with property variable z
Properties
{x} property variable x
KEY(x), VAL(x) key and value of property variable x
Labels
|x| label set variable x
c ELEMENTOF x check if label c exists in label set variable x
@@ 3.1 Working with Data, Labels, and Properties
In general, this capability of MetaPGML to bind property names and label sets to variables allows the seamless promotion of data to metadata, and vice versa, as is critical in applications such as data integration, cleaning, and governance. The $Q_{3}$ and $Q_{4}$ examples demonstrate such promotions and demotions of (meta)data. The following example aims to match the relationships between publications and the indexing databases and return the results in a structure to show in which indexing database each publication is being archived or indexed. $Q_{3}$: Which relationships does each publication have with each indexing databases? Query results are shown below, where LABEL resolves label sets bound to y, while z.Name values form column headers: (for convenience we switch between tabular and binding set representations of query results) Result bindings of $Q_{3}$ $\{$ $\left(\ \text{Title}\mapsto{\color[rgb]{0,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0.5}\text{"Nature Studies"}},\text{Scopus}\mapsto{\color[rgb]{0,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0.5}\text{\{"Archived"\}}}\ \right),$ $\left(\ \text{Title}\mapsto{\color[rgb]{0,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0.5}\text{"Nature Studies"}},\text{PubMed}\mapsto{\color[rgb]{0,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0.5}\text{\{"Indexed"\}}}\ \right),$ $\left(\ \text{Title}\mapsto{\color[rgb]{0,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0.5}\text{"Biology Advancements"}},\text{PubMed}\mapsto{\color[rgb]{0,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0.5}\text{\{"Indexed"\}}}\ \right)\}$ Our next query ($Q_{4}$) demonstrates how the MetaGPML can manipulate metadata as regular data. This query finds potential reviewers for publications based on their research fields and the publication’s acceptance criteria, showcasing how MPG treats metadata (property keys) as queryable data, enabling more flexible and powerful queries than the PG model. $Q_{4}$: Who are potential reviewer candidates for each publication based on the candidate’s research fields? Result table of $Q_{4}$
Reviewer candidate Publication venue Research field
Lee Nature Studies Biology
Lee Biology advancement Biology
Rose Nature Studies Ecology
The structure of the graph requires demoting the metadata of the property key in Publication, indicating when it started publishing in a certain research field. This allows comparison with the ResearchField property of Person entities, enabling identification of reviewer candidates for different publications.
@@ 3.2 Working with Reification
MPG and MetaGPML enable reification as the second type of metadata on property graphs. A sub-structure of an MPG can be reified as a node, thereby making it a first-class citizen. This enables annotating part of the graph by assigning properties and labels to a node that reifies that part. It is worth mentioning that the part of the graph that is being reified may or may not be a proper graph, which means that a node can reify some data objects without their assigned relationships in the graph, e.g., a set of specific properties or relationships, can be reified without the nodes they are associated with.
In Figure 1, we represent the statement “Rose assigned Lee as a reviewer on 5th November 2024” using node $n_{4}$ that induces a sub-structure containing node $n_{1}$, some relevant properties, and outgoing review edges. Query $Q_{5}$ retrieves the name of the person who assigns a reviewer:
@@
While it may look that MPG is a hypergraph model it’s important to note that hypergraphs focus on complex relationships, while reification focuses on complex objects. In MPG, edges still connect exactly two nodes. Nodes can act as meta-nodes, with other data objects (nodes, edges, properties, label sets) assigned for reification.
@@ 4 Meta-Property Graph Model
  • $\lambda:N\cup E\rightarrow L$ is a bijective labeling function assigning a label set identifier to each node and edge,
@@
where $E=E_{d}\cup E_{u}$ the set of compatible property sets $\mathbb{C}=\{C\subseteq P\mid\forall p_{1},p_{2}\in C,p_{1}\neq p_{2}\Rightarrow Key(p_{1})\neq Key(p_{2})\}$, $Key(p)=\pi_{1}(\upsilon(p))$, $Val(p)=\pi_{2}(\upsilon(p))$.
@@
The constraints we have imposed in our design of MPGs are very minimalistic, ensuring only that all properties are uniquely assigned to edges and nodes, and that reification is well-founded. We have deliberately avoided imposing unwarranted constraints. For instance, there is no constraint preventing a node from having an outgoing edge to a subgraph containing one of the node’s own properties. This flexibility allows for representing scenarios such as “Mary legally changed her first name” or “Mary audits all names in the database (including her own)”. Similarly, there is no constraint requiring that for each edge in $\rho(n)$, the source and target of the edge must also be in $\rho(n)$, which, for instance, can be particularly valuable in applications with privacy preservation concerns. As an example, This permits representations like “Mary is confident that Bob edited a book, but not confident about which book”, where only the node representing Bob and Bob’s outgoing edge labeled ’edited’ are included, without the target of this edge. Of course, such additional constraints can be added (or existing constraints can be removed) as appropriate to the specific application domain.
@@ 5 Meta-Property Graph Pattern Matching Language
We next formalize our Graph Pattern Matching Language for Metaproperty Graphs (MetaGPML) as an extension to the Graph Pattern Matching Language (GPML) [deutsch_graph_2022, francis_gpc_2023, francis_researchers_2023]. GPML is the core element of both SQL/PGQ and GQL languages standardized by ISO. By extending GPML we ensure backward compatibility in the sense that every GPML query is a MetaGPML query.
@@ 5.1 Syntax
Let $\mathcal{X}$ be a countable infinite set of variables. We introduce descriptors $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ in MetaGPML, consistent with GPML [francis_gpc_2023]. For $\text{{x}},\text{{y}},\text{{z}}\in\mathcal{X},\ \ell\in\mathcal{L},\ \mathcal{k}\in\mathcal{K},$ and $\mathcal{v}\in\mathcal{V}$, we have
Descriptors: $\delta_{nodes}$ $=\delta_{edges}\mid\text{{::}}\pi\mid\text{{:}}\ell\text{{::}}\pi\mid\text{{x::}}\pi\mid\text{{x:}}\ell\text{{::}}\pi$ (1) $\quad\mid\text{{:?y::}}\pi\mid\text{{x:?y::}}\pi$ $\delta_{edges}$ $=\text{{x}}\mid\text{{:}}\ell\mid\text{{:?y}}\mid\text{{x:}}\ell\mid\text{{x:?y}}$
The descriptors in (1) will be used in the following patterns to assign variables and constants to different data objects. A pattern $\pi$ in MetaGPML and operations on it can be formulated as follows:
[…]
We define $\mathcal{E}$ as the expression for condition $\Phi$ that can be used in a pattern after the WHERE keyword as follows, considering $c\in\mathcal{K}\cup\mathcal{V}\cup\mathcal{L}$.
[…]
It is important to note that $\text{{KEY}}(\text{{x}})$ and $\text{{VAL}}(\text{{x}})$ operate on properties as first-class data objects, similar to nodes and edges. The same applies to $\text{{LABEL}}(\text{{x}})$, which operates on label sets as distinguishable data objects and retrieves the set of label strings it contains. These functions should not be confused with similar functions in some other PG query languages like OpenCypher, which only simulates these capabilities at the query language level within the conventional PG data model. Finally, based on (15) and (21), we formulate the definition of clauses and queries.
[…]
@@ 5.2 Semantics
Next, we present the semantics of our language as a clear, unambiguous foundation for further study and implementation. To avoid unnecessary complexity, we presented the most vital semantics to illustrate the concept to understand how MetaGPML works. The complete and more detailed semantics covering all situations are presented in appendix A.
@@
(Bindings) A binding $\beta:\mathcal{X}\rightarrow\mathbb{V}$ assigns variables $x\in\mathcal{X}$ to values $v\in\mathbb{V}$, where $\mathbb{V}=\mathcal{I}\cup\mathcal{L}\cup\mathcal{K}\cup\mathcal{V}$. We denote this as $(x_{1}\mapsto v_{1},...,x_{n}\mapsto v_{n})$, where $x_{1},...,x_{n}$ are the variables in the domain of $\beta$ $(\text{Dom}(\beta))$ and $v_{1},...,v_{n}$ are their corresponding values. We denote the empty binding as ()()\,.
@@
Node pattern matching: $\llbracket\text{{()}}\rrbracket_{G}=\{()\mid n\in N\}$ (27) $\llbracket\text{{(x)}}\rrbracket_{G}=\{(\text{{x}}\mapsto n)\mid n\in N\}$ (28) $\llbracket\text{{(x:$\ell$)}}\rrbracket_{G}=\{(\text{{x}}\mapsto n)\mid n\in N,\ell\in\mu(\lambda(n))\}$ (29) $\llbracket\text{{(x:?y)}}\rrbracket_{G}=\{(\text{{x}}\mapsto n,\text{{y}}\mapsto l)\mid n\in N,l=\lambda(n)\}$ (30) $\llbracket\text{{(x).z}}\rrbracket_{G}=\{(\text{{x}}\mapsto n,\text{{z}}\mapsto p)\mid n\in N,p\in\sigma(n)\}$ (31)
@@
Meta-nodes are a key concept in MetaGPML, defined as nodes to which diverse elements are assigned based on the concept of sub-structure. This feature enables reification within the Meta-Property Graph, allowing for a more nuanced representation of complex relationships and structures.
@@
Meta-node pattern matching: $\llbracket\text{{(x:$\ell$::}}\pi\text{{)}}\rrbracket_{G}=\llbracket\text{{(x:$\ell$)}}\rrbracket_{G}\Join\llbracket\pi\rrbracket_{G_{n}}$ (32)
The meta-node semantic definition involves a full join operation between two patterns divided by :: in the node pattern descriptor. Here, $\llbracket\text{{(x:$\ell$)}}\rrbracket_{G}$ represents a simple node pattern matching, while $\llbracket\pi\rrbracket_{G_{n}}$ is a pattern matching within the $G_{n}$ sub-structure. We define the semantics of edges in our model, supporting both directed and undirected edges.
Edge pattern matching: -$\llbracket\text{{-[]->}}\rrbracket_{G}$ $=\{()\mid e\in E_{d}\}$ (33) -$\llbracket\text{{-[x]->}}\rrbracket_{G}$ $=\{(\text{{x}}\mapsto e)\mid e\in E_{d}\}$ (34) -$\llbracket\text{{-[:$\ell$]->}}\rrbracket_{G}$ $=\{()\mid e\in E_{d},\ell\in\mu(\lambda(e))\}$ (35) -$\llbracket\text{{-[x:$\ell$]-}}\rrbracket_{G}$ $=\{(\text{{x}}\mapsto e)\mid e\in E_{u},\ell\in\mu(\lambda(e))\}$ (36) -$\llbracket\text{{-[x:?y]->}}\rrbracket_{G}$ $=\{(\text{{x}}\mapsto e,\text{{y}}\mapsto l)\mid e\in E_{d},l=\lambda(e)\}$ (37) -$\llbracket\text{{-[x].z->}}\rrbracket_{G}$ $=\{(\text{{x}}\mapsto e,\text{{z}}\mapsto p)\mid e\in E_{d},p\in\sigma(e)\}$ (38)
Equations 33-81 match various edge patterns, including directed and undirected edges with or without labels. Equation 37 allows assigning the edge’s label set to variable y, while 38 enables assigning the edge’s properties to variable z.
@@
Property and label objects: $\llbracket\text{{\{x\}}}\rrbracket_{G}$ $=\{(\text{{x}}\mapsto p)\mid p\in P\}$ (39) $\llbracket\text{{|x|}}\rrbracket_{G}$ $=\{(\text{{x}}\mapsto l)\mid l\in L\}$ (40)
[…]
In (42), $\beta^{\prime}$ maps any variable in ${\llbracket\pi_{1}\text{{+}}\pi_{2}\rrbracket}_{G}$ not in $\beta$’s domain to $Null$.
@@
The semantics $\llbracket\mathcal{E}\rrbracket^{\beta}_{G}$ of an expression $\mathcal{E}$ is computed with respect to binding $\beta$ over graph $G$. For a variable x in $\beta$’s domain, $\llbracket\text{{x}}\rrbracket^{\beta}_{G}=\beta(x)$.
Node and property values: $\llbracket\text{{x.}}\mathcal{k}\rrbracket^{\beta}_{G}=\begin{cases}Val(p)&\text{if }\beta(\text{{x}})\in N\cup E,p\in\sigma(\beta(\text{{x}})),Key(p)=\mathcal{k}\\ Null&\text{otherwise}\end{cases}$ (45)
Property and Label set operations: $\llbracket\text{{KEY(x)}}\rrbracket^{\beta}_{G}=Key(\llbracket\text{{x}}\rrbracket^{\beta}_{G})\text{ for }\llbracket\text{{x}}\rrbracket^{\beta}_{G}\in P$ (46) $\llbracket\text{{VAL(x)}}\rrbracket^{\beta}_{G}=Val(\llbracket\text{{x}}\rrbracket^{\beta}_{G})\text{ for }\llbracket\text{{x}}\rrbracket^{\beta}_{G}\in P$ (47) $\llbracket\text{{LABEL(x)}}\rrbracket^{\beta}_{G}=\mu(\llbracket\text{{x}}\rrbracket^{\beta}_{G})\text{ for }\llbracket\text{{x}}\rrbracket^{\beta}_{G}\in L$ (48)